# On decoding extended Han-Zhang codes

### v1

@@ 26          ${\bf c}\leftarrow{\bf y}-{\bf x}_{0}$
  • Note that $\mathcal{S}$} contains a $7$-zero-sum subset, such as the $7$-subset $\{1,{\omega}^{12},{\omega}^{2},{\omega}^{13},{\omega}^{3},{\omega}^{8},{\omega}^{9}\}$. It then follows from Lemmas 3 and 4.2) that the extended Han-Zhang code ${\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)$ is a $14,7,7]_{16}$ NMDS code with parity-check matrix
    $H_{{\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)}=\left(\begin{array}[]{cccccccccccccc}1&{\omega}^{11}&{\omega}^{7}&{\omega}^{11}&{\omega}^{14}&{\omega}^{7}&{\omega}^{11}&{\omega}^{13}&{\omega}^{13}&{\omega}^{13}&{\omega}^{14}&{\omega}^{14}&{\omega}^{7}&0\\ 1&{\omega}^{12}&{\omega}^{9}&{\omega}^{14}&{\omega}^{3}&{\omega}^{13}&{\omega}^{3}&{\omega}^{6}&{\omega}^{7}&{\omega}^{9}&{\omega}^{11}&{\omega}^{12}&{\omega}^{6}&0\\ 1&{\omega}^{13}&{\omega}^{11}&{\omega}^{2}&{\omega}^{7}&{\omega}^{4}&{\omega}^{10}&{\omega}^{14}&{\omega}&{\omega}^{5}&{\omega}^{8}&{\omega}^{10}&{\omega}^{5}&0\\ 1&{\omega}^{14}&{\omega}^{13}&{\omega}^{5}&{\omega}^{11}&{\omega}^{10}&{\omega}^{2}&{\omega}^{7}&{\omega}^{10}&{\omega}&{\omega}^{5}&{\omega}^{8}&{\omega}^{4}&0\\ 1&1&1&{\omega}^{8}&1&{\omega}&{\omega}^{9}&1&{\omega}^{4}&{\omega}^{12}&{\omega}^{2}&{\omega}^{6}&{\omega}^{3}&0\\ 1&{\omega}&{\omega}^{2}&{\omega}^{11}&{\omega}^{4}&{\omega}^{7}&{\omega}&{\omega}^{8}&{\omega}^{13}&{\omega}^{8}&{\omega}^{14}&{\omega}^{4}&{\omega}^{2}&1\\ 1&{\omega}^{2}&{\omega}^{4}&{\omega}^{14}&{\omega}^{8}&{\omega}^{13}&{\omega}^{8}&{\omega}&{\omega}^{7}&{\omega}^{4}&{\omega}^{11}&{\omega}^{2}&{\omega}&1\end{array}\right).$
@@
  • Modifying Lines 8 to 14 of Algorithm 1 based on Table I, we have $\ell=3$ and ${{\mathbf{u}}}=(1,{\omega}^{11},{\omega}^{7},{\omega}^{11},{\omega}^{14},{\omega}^{7},{\omega}^{11},{\omega}^{13},{\omega}^{13},{\omega}^{13},\\ {\omega}^{14},{\omega}^{14},{\omega}^{7})$. Then $(\mathcal{A},\mathcal{B})$ is a $3$-ECP of ${\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)$, where
    $\mathcal{A}=\{(f(a_{1}),\ldots,f(a_{n}),0):~f(x)\in\mathbb{F}_{q}[x]_{4}\}~\mbox{and}~\mathcal{B}={\mathrm{GRS}}_{3}({\mathcal{S}},{\bf v}^{-1}\star{\bf u},\infty).$
    Moreover, we deduce that
    $G_{\mathcal{A}}=\left(\begin{array}[]{cccccccccccccc}1&1&1&1&1&1&1&1&1&1&1&1&1&0\\ 1&{\omega}&{\omega}^{2}&{\omega}^{3}&{\omega}^{4}&{\omega}^{6}&{\omega}^{7}&{\omega}^{8}&{\omega}^{9}&{\omega}^{11}&{\omega}^{12}&{\omega}^{13}&{\omega}^{14}&0\\ 1&{\omega}^{2}&{\omega}^{4}&{\omega}^{6}&{\omega}^{8}&{\omega}^{12}&{\omega}^{14}&{\omega}&{\omega}^{3}&{\omega}^{7}&{\omega}^{9}&{\omega}^{11}&{\omega}^{13}&0\\ 1&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&0\\ \end{array}\right)~\mbox{and}$~
    $G_{\mathcal{B}}=\left(\begin{array}[]{cccccccccccccc}1&1&1&1&1&1&1&1&1&1&1&1&1&0\\ 1&{\omega}^{12}&{\omega}^{9}&{\omega}^{14}&{\omega}^{3}&{\omega}^{13}&{\omega}^{3}&{\omega}^{6}&{\omega}^{7}&{\omega}^{9}&{\omega}^{11}&{\omega}^{12}&{\omega}^{6}&0\\ 1&{\omega}^{13}&{\omega}^{11}&{\omega}^{2}&{\omega}^{7}&{\omega}^{4}&{\omega}^{10}&{\omega}^{14}&{\omega}&{\omega}^{5}&{\omega}^{8}&{\omega}^{10}&{\omega}^{5}&1\\ \end{array}\right).$
@@
    ${\bf a}={\bf s}_{0}G_{\mathcal{A}}=({\omega}^{5},{\omega}^{12},{\omega}^{13},{\omega}^{13},{\omega},0,{\omega}^{5},{\omega}^{4},{\omega}^{10},{\omega}^{8},{\omega}^{13},{\omega},{\omega}^{5},0)~{\rm and}~Z=z({\bf a})=\{6,14\}.$
@@
IV Deep holes of extended Han-Zhang codes
In addition to the existence of ECPs, deep holes of linear codes are also closely related to their maximum-likelihood decoding problems. In this section, we discuss the deep holes of the extended Han-Zhang codes and apply the deep holes to obtain more non-GRS MDS codes with larger lengths and dimensions. Moreover, we discuss their connections to the well-known Roth-Lempel codes. According to Definition 6, we first need to determine the covering radius of extended Han-Zhang codes.
Theorem 16.
For any extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$, we have
$\rho({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=n-k+1.$
Proof.
With Definition 4, it follows from the redundancy bound ( [21, Corollary 11.1.3]) that
$\rho({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))\leq n-k+1.$ (19)
On the other hand, by Lemma 12, we have that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)\subseteq{\mathrm{GRS}}_{k+1}({\mathcal{S}},{\bf v},\infty)$, where ${\mathrm{GRS}}_{k+1}({\mathcal{S}},{\bf v},\infty)$ is an $n+1,k+1,n-k+1]_{q}$ MDS code. Combining with the supercode lemma ( [21, Corollary 11.1.5]), it then implies that
$\begin{split}\rho({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))\geq\min\{{{\rm{wt}}}(\mathbf{c}):\mathbf{c}\in{\mathrm{GRS}}_{k+1}({\mathcal{S}},{\bf v},\infty)\setminus{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)\}\geq n-k+1.\end{split}$ (20)
We complete the proof by considering (19) and (20) together. ∎
IV-A Determining deep holes of extended Han-Zhang codes
Zhuang etal.et~al. provided in [58, Proposition 2] an equivalent condition under which a vector is a deep hole of a Reed-Solomon code, utilizing the covering radius and a generator matrix of this code. More recently, Wu etal.et~al. extended this result to general MDS codes in [48, Lemma 5]. However, as shown in Lemma 4, the code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is not always an MDS code. Consequently, the results from [48, Lemma 5] and [58, Proposition 2] cannot be directly applied to determine the deep holes of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. This limitation motivates the following theorem.
Theorem 17.
Suppose that $\mathbf{u}\in\mathbb{F}_{q}^{n+1}$ and $\left(\begin{array}[]{c}G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}\\ \mathbf{u}\end{array}\right)$ generates a linear code $\mathcal{C}$$^{\prime}$. Then $\mathbf{u}$ is a deep hole of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ if and only if $\mathcal{C}$$^{\prime}$ is an $n+1,k+1,n-k+1]_{q}$ MDS code.
Proof.
Note that if $\mathbf{u}$ is a deep hole of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$, then $\mathbf{u}\notin{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$, which implies that $\mathcal{C}$$^{\prime}$ is an $n+1,k+1]_{q}$ linear code. Note that
$d({\mathcal{C}}^{\prime})=\min\{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)),d(\mathbf{u},{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))\}$
and it follows from Lemmas 4 and 16 as well as Definition 6 that
$d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))\geq n-k+1=\rho({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))\geq d(\mathbf{u},{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)).$
Hence, $d({\mathcal{C}}^{\prime})=d(\mathbf{u},{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))$. It further implies that $\mathbf{u}$ is a deep hole of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ if and only if $d({\mathcal{C}}^{\prime})=d(\mathbf{u},{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=\rho({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=n-k+1$, if and only if $\mathcal{C}$$^{\prime}$ is an $n+1,k+1,n-k+1]_{q}$ MDS code. This completes the proof. ∎
Lemma 18.
Suppose that $\mathbf{g}=(g(a_{1}),\ldots,g(a_{n}),u_{n+1})\in\mathbb{F}_{q}^{n+1}$, where $g(x)=g^{\prime}(x)+f(x)\in\mathbb{F}_{q}[x]$ with $f(x)\in\mathcal{V}_{k}$. If $(\mathbf{v},v_{n+1})\star\mathbf{g}$ is a deep hole of the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ for $v_{n+1}\in\mathbb{F}_{q}$, then $g^{\prime}(x)\in\mathbb{F}_{q}[x]\setminus\mathcal{V}_{k}$.
Proof.
Assume that $g^{\prime}(x)\in\mathcal{V}_{k}$. Then $g(x)=g^{\prime}(x)+f(x)\in\mathcal{V}_{k}$ and
$(\mathbf{v},v_{n+1})\star\mathbf{g}=(v_{1}(g^{\prime}(a_{1})+f(a_{1})),\ldots,v_{n}(g^{\prime}(a_{n})+f(a_{n})),u_{n+1}v_{n+1}).$
It can be checked that
$\left(\begin{array}[]{c}G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}\\ (\mathbf{v},v_{n+1})\star\mathbf{g}\end{array}\right)=\left(\begin{array}[]{ccccc}v_{1}&\ldots&v_{n}&0\\ v_{1}a_{1}&\ldots&v_{n}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{1}a_{1}^{k-2}&\ldots&v_{n}a_{n}^{k-2}&0\\ v_{1}a_{1}^{k}&\ldots&v_{n}a_{n}^{k}&1\\ v_{1}(g^{\prime}(a_{1})+f(a_{1}))&\ldots&v_{n}(g^{\prime}(a_{n})+f(a_{n}))&u_{n+1}v_{n+1}\end{array}\right)$
and it can be transformed to the matrix
$\left(\begin{array}[]{ccccc}v_{1}&\ldots&v_{n}&0\\ v_{1}a_{1}&\ldots&v_{n}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{1}a_{1}^{k-2}&\ldots&v_{n}a_{n}^{k-2}&0\\ v_{1}a_{1}^{k}&\ldots&v_{n}a_{n}^{k}&1\\ 0&\ldots&0&u_{n+1}v_{n+1}-f_{k}-g_{k}\end{array}\right)$ (27)
by a series of elementary row operations, where $f_{k}$ and $g_{k}$ are respective coefficients of $x^{k}$ in $f(x)$ and $g^{\prime}(x)$. From Definition 4, we have $k+1\leq n$. It then follows that the determinants of all the $(k+1)\times(k+1)$ matrices consisting of any $k+1$ columns of the first $n$ columns of the matrix given in (27) equal zero. With Lemma 10, we immediately conclude that the linear code generated by $\left(\begin{array}[]{c}G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}\\ (\mathbf{v},v_{n+1})\star\mathbf{g}\end{array}\right)$ is not MDS. According to Theorem 17, we get that $g^{\prime}(x)\in\mathbb{F}_{q}[x]\setminus\mathcal{V}_{k}$ if $(\mathbf{v},v_{n+1})\star\mathbf{g}$ is a deep hole of the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. This completes the proof. ∎
With Lemma 18, we can always assume that no term in $g^{\prime}(x)$ belongs to $\mathcal{V}_{k}$. In the following, we consider two cases of $g^{\prime}(x)$ in terms of their lowest degrees. In other words, we consider $g^{\prime}(x)=g_{k-1}x^{k-1}$ with $g_{k-1}\in\mathbb{F}_{q}^{*}$ and $g^{\prime}(x)=g_{k+1}x^{k+1}+g_{k-1}x^{k-1}$ with $g_{k+1}\in\mathbb{F}_{q}^{*}$ and $g_{k-1}\in\mathbb{F}_{q}$ in the following.
Theorem 19.
Suppose that $\mathbf{g}=(g(a_{1}),\ldots,g(a_{n}),u_{n+1})\in\mathbb{F}_{q}^{n+1}$, where $g(x)=g_{k-1}x^{k-1}+f(x)\in\mathbb{F}_{q}[x]$ with $g_{k-1}\in\mathbb{F}_{q}^{*}$ and $f(x)\in\mathcal{V}_{k}$. Then for any $v_{n+1}\in\mathbb{F}_{q}$, $(\mathbf{v},v_{n+1})\star\mathbf{g}$ is a deep hole of the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ if and only if one of the following conditions holds:
@@ IV-A Determining deep holes of extended Han-Zhang codes
    $u_{n+1}v_{n+1}=f_{k}$;

### v2

@@ IV-C Decoding examples of success and failure
  • Note that $\mathcal{S}$} contains a $7$-zero-sum subset, such as the $7$-subset $\{1,{\omega}^{12},{\omega}^{2},{\omega}^{13},{\omega}^{3},{\omega}^{8},{\omega}^{9}\}$. It then follows from Lemmas 2 and 3.2) that the ESGRS code ${\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)$ is a $14,7,7]_{16}$ NMDS code with parity-check matrix of the form (13).
    $H_{{\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)}=\left(\begin{array}[]{cccccccccccccc}1&{\omega}^{11}&{\omega}^{7}&{\omega}^{11}&{\omega}^{14}&{\omega}^{7}&{\omega}^{11}&{\omega}^{13}&{\omega}^{13}&{\omega}^{13}&{\omega}^{14}&{\omega}^{14}&{\omega}^{7}&0\\ 1&{\omega}^{12}&{\omega}^{9}&{\omega}^{14}&{\omega}^{3}&{\omega}^{13}&{\omega}^{3}&{\omega}^{6}&{\omega}^{7}&{\omega}^{9}&{\omega}^{11}&{\omega}^{12}&{\omega}^{6}&0\\ 1&{\omega}^{13}&{\omega}^{11}&{\omega}^{2}&{\omega}^{7}&{\omega}^{4}&{\omega}^{10}&{\omega}^{14}&{\omega}&{\omega}^{5}&{\omega}^{8}&{\omega}^{10}&{\omega}^{5}&0\\ 1&{\omega}^{14}&{\omega}^{13}&{\omega}^{5}&{\omega}^{11}&{\omega}^{10}&{\omega}^{2}&{\omega}^{7}&{\omega}^{10}&{\omega}&{\omega}^{5}&{\omega}^{8}&{\omega}^{4}&0\\ 1&1&1&{\omega}^{8}&1&{\omega}&{\omega}^{9}&1&{\omega}^{4}&{\omega}^{12}&{\omega}^{2}&{\omega}^{6}&{\omega}^{3}&0\\ 1&{\omega}&{\omega}^{2}&{\omega}^{11}&{\omega}^{4}&{\omega}^{7}&{\omega}&{\omega}^{8}&{\omega}^{13}&{\omega}^{8}&{\omega}^{14}&{\omega}^{4}&{\omega}^{2}&1\\ 1&{\omega}^{2}&{\omega}^{4}&{\omega}^{14}&{\omega}^{8}&{\omega}^{13}&{\omega}^{8}&{\omega}&{\omega}^{7}&{\omega}^{4}&{\omega}^{11}&{\omega}^{2}&{\omega}&1\end{array}\right)$ (13)
@@
  • Modifying Lines 8 to 14 of Algorithm 1 based on Table I, we have $\ell=3$ and ${{\mathbf{u}}}=(1,{\omega}^{11},{\omega}^{7},{\omega}^{11},{\omega}^{14},{\omega}^{7},{\omega}^{11},{\omega}^{13},\\ {\omega}^{13},{\omega}^{13},{\omega}^{14},{\omega}^{14},{\omega}^{7})$. Then $(\mathcal{A},\mathcal{B})$ is a $3$-ECP of ${\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)$, where
    $\mathcal{A}=\{(f(a_{1}),\ldots,f(a_{n}),0):~f(x)\in\mathbb{F}_{q}[x]_{4}\}$
    and
    $\mathcal{B}={\mathrm{GRS}}_{3}({\mathcal{S}},{\bf v}^{-1}\star{\bf u},\infty).$
    Moreover, we deduce (14) and (15).
    $G_{\mathcal{A}}=\left(\begin{array}[]{cccccccccccccc}1&1&1&1&1&1&1&1&1&1&1&1&1&0\\ 1&{\omega}&{\omega}^{2}&{\omega}^{3}&{\omega}^{4}&{\omega}^{6}&{\omega}^{7}&{\omega}^{8}&{\omega}^{9}&{\omega}^{11}&{\omega}^{12}&{\omega}^{13}&{\omega}^{14}&0\\ 1&{\omega}^{2}&{\omega}^{4}&{\omega}^{6}&{\omega}^{8}&{\omega}^{12}&{\omega}^{14}&{\omega}&{\omega}^{3}&{\omega}^{7}&{\omega}^{9}&{\omega}^{11}&{\omega}^{13}&0\\ 1&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&0\\ \end{array}\right)$ (14)
    $G_{\mathcal{B}}=\left(\begin{array}[]{cccccccccccccc}1&1&1&1&1&1&1&1&1&1&1&1&1&0\\ 1&{\omega}^{12}&{\omega}^{9}&{\omega}^{14}&{\omega}^{3}&{\omega}^{13}&{\omega}^{3}&{\omega}^{6}&{\omega}^{7}&{\omega}^{9}&{\omega}^{11}&{\omega}^{12}&{\omega}^{6}&0\\ 1&{\omega}^{13}&{\omega}^{11}&{\omega}^{2}&{\omega}^{7}&{\omega}^{4}&{\omega}^{10}&{\omega}^{14}&{\omega}&{\omega}^{5}&{\omega}^{8}&{\omega}^{10}&{\omega}^{5}&1\\ \end{array}\right)$ (15)
@@
    ${\bf a}$ $={\bf s}_{0}G_{\mathcal{A}}$
    $=({\omega}^{5},{\omega}^{12},{\omega}^{13},{\omega}^{13},{\omega},0,{\omega}^{5},{\omega}^{4},{\omega}^{10},{\omega}^{8},{\omega}^{13},{\omega},{\omega}^{5},0)$
    and
    $Z=z({\bf a})=\{6,14\}.$
@@
V Concluding remarks
ESGRS codes ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ constitute a class of linear codes where each code is either a non-GRS MDS code or an NMDS code, both of which have significant applications in distributed storage systems and cryptography. Although various properties of these codes, including their dual codes, Schur squares, weight distributions, and (almost) self-dual properties, were investigated in [25, 1], their decoding challenges have remained largely unexplored.
This paper aims to bridge this gap, i.e., to answer Problem 1. We fully examined the existence of $\ell$-ECPs $(\mathcal{A},\mathcal{B})$ for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ and provided explicit constructions for $\mathcal{A}$ and $\mathcal{B}$, either via subcode relationships or through direct constructions, as established in Theorems 9 and 10. Note that these ECPs are composed of GRS codes and their variants, and are always defined over the base field. Based on these $\ell$-ECPs, we proposed a decoding algorithm for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ that can uniquely correct up to $\ell$ errors in $O(n^{3})$ time, as detailed in Theorem 11 and Algorithm 1. As discussed in Subsection IV-B, our decoding algorithm is more efficient than some alternative approaches.
We also highlight several research directions that warrant further exploration, summarized as follows.
@@ V Concluding remarks
    In Subsection IV-B, we mentioned two potential approaches to further reduce the decoding complexity of Algorithm 1. Therefore, it would be interesting to design $\mathcal{S}$} and $\bf v$} such that the above two approaches can be reached for some special error patterns.